Write a quadratic equation with -3/4 and 4 as its roots. Wite the equation in form ax^2 + bx + c = 0, where a,b and c are intergers.
Answer by ankor@dixie-net.com(22740)

Write a quadratic equation with -3/4 and 4 as its roots.
:
These two roots could be derived from the factors:
(4x + 3) (x - 4) = 0
FOIL
4x^2 - 16x + 3x - 12 = 0
Combine like terms
4x^2 - 13x - 12 = 0